OpenStudy (anonymous):

the turn signal on car A blinks 7 times every 5 seconds the turn signal on car B blinks 5 times every 4 seconds in one minute how many more times will the turn signal blink on car A than the turn signal on car B a.9 b.15 c.75 or d.84

OpenStudy (anonymous):

a.9 total=60secs/5=12, 12*7=84 60/4=15,15*5=75 84-75=9